OpenStudy (karatechopper):

write an equation in slope intercept form for the line -parallel to 2x+y=5- n with given conditions of point (6,2)

OpenStudy (angela210793):

parallel--->slope is the same=-2 y=5-n-2x y-2=-2(x-6) y=-2x+12+2 y=-2x+14

OpenStudy (karatechopper):

n was not part of the question

OpenStudy (angela210793):

ahhha..ok it doesn't change anything :)

OpenStudy (karatechopper):

i still dont get it

OpenStudy (anonymous):

2x+ y = 5 y = -2x + 5 m = - 2 y = - 2x +c (6,2) 2 = -12 + c c = 14 y = -2x + 14

OpenStudy (karatechopper):

should the c not be 5?

OpenStudy (angela210793):

no cause 12+2=14 not 5